OpenStudy (otonogold):

Check my answer? Use natural logarithms to solve the equation. round to the nearest thousandth. 3e^2x+4=24 I got 0.949 as my answer.

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large 3e^{2x+4}=24\] ??

OpenStudy (otonogold):

yes

jimthompson5910 (jim_thompson5910):

That is incorrect. The answer should be negative.

OpenStudy (otonogold):

Is the equation \[\Large 3e^{2x+4}=24\] this is went I saw you type. so 3e^2x plus 4 = 24 would be negative?

OpenStudy (otonogold):

this is what I saw **

jimthompson5910 (jim_thompson5910):

IF the equation is \[\Large 3e^{2x} + 4 = 24\] then the answer is roughly \(\Large x \approx 0.949\)

jimthompson5910 (jim_thompson5910):

Notice how in this new equation the +4 is not part of the exponent

OpenStudy (otonogold):

Lol Im so upset I cant see the equation you did :( so it's really hard to understand :P

jimthompson5910 (jim_thompson5910):

Oh you can't see the equations I'm posting? Ok I'll draw them out

jimthompson5910 (jim_thompson5910):

Which is the original equation?